OpenStudy (anonymous):

the perimeter of a rectangle is twice the sum of its length and its width. the perimeter is 22 meters and its lenght is 2 meters more than twice its width, what is the length?

OpenStudy (anonymous):

p= 2(l+w)=22 l=2+2w ---> w= (l-2)/2 2(l+(l-2)/2)=22 2l+l-2=22 3l=24 l=8 (w=3)